OpenStudy (anonymous):

Which ordered pair is in the solution set of the system of equations y=-x+1 and y=x^2+5x+6? a) (-5,-1) b) (-5,6) c) (5,-4) d) (5,2)

myininaya (myininaya):

-x+1=x^2+5x+6 x^2+6x+5=0 (x+1)(x+5)=0 x=-1 x=-5 if x=-5, then using one of the eqations for y (doesn't matter which) we can find the corresponding y value for x=-5. y=-(-5)+1=6 B

OpenStudy (anonymous):

i dont get what the question is asking lol.